OpenStudy (anonymous):

24x^3/25y^2 *10y^5/18x

OpenStudy (anonymous):

\[\frac{ 24x^3 }{ 25y^2 } * \frac{ 10y^5 }{ 18x }\]

OpenStudy (anonymous):

You got it! Now what do you do next?

OpenStudy (anonymous):

I have no idea, thats what i need help with.

OpenStudy (anonymous):

Try cross multiply, let's see the results. If that doesn't work then multiply straight horizontally.

OpenStudy (anonymous):

\[\frac{ 450x^3 }{ 250y^7 }\] thats what i got for cross multiply

OpenStudy (anonymous):

just kidding top was 432x^3

OpenStudy (anonymous):

@CGGURUMANJUNATH, could you help this woman? 'Cause at the same time I'm working with someone else at the moment.

OpenStudy (anonymous):

thank you

OpenStudy (cggurumanjunath):